# Explanation as a Watermark: Towards Harmless and Multi-bit Model Ownership Verification via Watermarking Feature Attribution

### v1

@@ V-C2 The Resistance to Model-pruning Attack
Model-pruning serves as a potential watermark-removal attack because it may prune watermark-related neurons. We exploit parameter pruning [16] as an example for discussion. Specifically, we prune the neurons in the model by zeroing out those with the lowest $l_{1}$ norm. In particular, we use pruning rate to denote how many proportions of neuron are pruned.
As shown in Figure 6, the test accuracy of ResNet-18 drops while the PPL of GPT-2 increases, as the pruning rate increases. It indicates the degradation of the functionality of the model. However, the p-value of the pruned model negligibly changes and is lower than the significant level $\alpha$. Besides, the WSR is still greater than $0.9$. These results suggest that our EaaW resists to model-pruning attack.
@@ V-C3 The Resistance to Adaptive Attacks
In practice, the adversaries may know the existence of our EaaW and design adaptive attacks to circumvent it. Specifically, they may try to remove the watermark or interfere with the watermark extraction by manipulating the explanation of the input data. Existing techniques for manipulating explanations can be classified into two categories: (1) modifying model parameters [19, 41] or (2) modifying the inputs ($i.e.$, the adversarial attack against XAI) [12]. In this section, we hereby present the results under the first type of attacks in two representative scenarios, namely the overwriting attack and the unlearning attack. The discussion about the second one can be found in Appendix -E.
@@
The experimental results of the overwriting attack are shown in Table V. We conduct a $10$-epoch fine-tuning to simulate the overwriting process and the models have already converged after 10 epochs. After the overwriting attack, the functionality of both the watermarked ResNet-18 and the watermarked GPT-2 decreases, while the p-values are still low and the WSRs are close to $0.9$. It indicates that the overwriting attack cannot effectively remove our EaaW watermark.
TABLE V: Watermark success rate (WSR), the log p-value, and functionality evaluation (test accuracy or PPL) of ResNet-18 and GPT-2 against overwriting attack and unlearning attack.
Model Metric $\downarrow$ Before After Overwriting After Unlearning
@@
The experimental results of the unlearning attack are illustrated in Table V. The results demonstrate that our EaaW also resists to unlearning attack. Specifically, the WSRs drop only $0.112$ and $0.031$, respectively. The watermark can still be extracted from the model and the ownership can still be verified with low p-values.
@@ VI-A How EaaW Affect the Watermarked Model?
From Figure 10, we can see that the intermediate feature representations are generally unaffected before and after embedding the watermark using EaaW. The feature of the trigger sample continues to reside within the cluster corresponding to its respective class (the red class numbered 00), thereby demonstrating the harmlessness of EaaW. Additionally, as depicted in Figure 10, EaaW actually changes the direction of the trigger sample to the decision boundary of the model, that is, the direction of the trigger sample’s gradient. By embedding a multi-bit watermark into the sign of this gradient, EaaW achieves successful model watermarking without altering the prediction of the trigger sample.
@@ VI-B Security Analysis against Ambiguity Attack
When the adversary acquires the watermarked model, the adversary can attempt to forge a fake watermark to establish its ownership of the watermarked model. If both the true model owner and the adversary can independently authenticate their copyright claims, it becomes impossible to ascertain the actual ownership. This type of attack is named ambiguity attack [11] (or false claim attack [32]). The definition of the ambiguity attack for our EaaW is as follows.

### v2

@@ V-C2 The Resistance to Model-pruning Attack
Model-pruning serves as a potential watermark-removal attack because it may prune watermark-related neurons. We exploit parameter pruning [17] as an example for discussion. Specifically, we prune the neurons in the model by zeroing out those with the lowest $l_{1}$ norm. In particular, we use the pruning rate to denote how many proportions of neurons are pruned.
As shown in Figure 5, the test accuracy of ResNet-18 drops while the PPL of GPT-2 increases, as the pruning rate increases. It indicates the degradation of the functionality of the model. However, the p-value of the pruned model negligibly changes and is lower than the significant level $\alpha$. Besides, the WSR is still greater than $0.9$. These results suggest that our EaaW resists the model-pruning attack.
@@ V-C3 The Resistance to Adaptive Attacks
In practice, the adversaries may know the existence of our EaaW and design adaptive attacks to circumvent it. Specifically, they may try to remove the watermark or interfere with the watermark extraction by manipulating the explanation of the input data. Existing techniques for manipulating explanations can be classified into two categories: (1) modifying model parameters [20, 44] or (2) modifying the inputs ($i.e.$, the adversarial attack against XAI) [13]. In this section, we hereby present the results under the first type of attacks in two representative scenarios, namely the overwriting attack and the unlearning attack. The discussion on the second attack can be found in Appendix -E.
@@
The experimental results of the overwriting attack are shown in Table V. We conduct a $10$-epoch fine-tuning to simulate the overwriting process and the models have already converged after 10 epochs. After the overwriting attack, the functionality of both the watermarked ResNet-18 and the watermarked GPT-2 decreases, while the p-values are still low and the WSRs of the original watermarks are close to $0.9$. It indicates that the overwriting attack cannot effectively remove our EaaW watermark. We notice that the overwriting attack can embed the adversary’s watermark into the victim model to some extent. As shown in Table V, the WSRs of the new watermarks are larger than $0.8$, indicating that there are two distinct watermarks within the model. However, this is a common and trivial situation in watermarking. This issue can be easily solved by registering the watermark and the model to a trusted third party ($e.g.$, the intellectual property office) accompanied by timestamps [59]. The watermark with a later timestamp will not be treated as a valid copyright certificate.
TABLE V: Watermark success rate (WSR) of the original watermark (dubbed ‘Ori. WM’) and the adversary’s new watermark (dubbed ‘New WM’), the log p-value, and functionality evaluation (test accuracy or PPL) of ResNet-18 and GPT-2 against overwriting attack and unlearning attack.
[…]
@@
The experimental results of the unlearning attack are illustrated in Table V. The results demonstrate that our EaaW also resists unlearning attacks. Specifically, the WSRs drop only $0.112$ and $0.031$, respectively. The watermark can still be extracted from the model and the ownership can still be verified with low p-values.
@@ VI-A How EaaW Affect the Watermarked Model?
From Figure 9, we can see that the intermediate feature representations are generally unaffected before and after embedding the watermark using EaaW. The feature of the trigger sample continues to reside within the cluster corresponding to its respective class (the red class numbered 00), thereby demonstrating the harmlessness of EaaW. Additionally, as depicted in Figure 9, EaaW actually changes the direction of the trigger sample to the decision boundary of the model, that is, the direction of the trigger sample’s gradient (as visualized as the black arrows towards the general decision boundary). By embedding a multi-bit watermark into the sign of this gradient, EaaW achieves successful model watermarking without altering the prediction of the trigger sample.
@@ VI-B Security Analysis against Ambiguity Attack
When the adversary acquires the watermarked model, the adversary can attempt to forge a fake watermark to establish its ownership of the watermarked model. If both the true model owner and the adversary can independently authenticate their copyright claims, it becomes impossible to ascertain the actual ownership. This type of attack is named ambiguity attack [12] or false claim attack [35]. More details are in Appendix -F. The ambiguity attack for our EaaW is as follows.
@@ VII Conclusion
Acknowledgment
This research is supported in part by the National Key Research and Development Program of China under Grant 2021YFB3100300 and the National Natural Science Foundation of China under Grants (62072395 and U20A20178). This work was mostly done when Yiming Li was at the State Key Laboratory of Blockchain and Data Security, Zhejiang University. He is currently at Nanyang Technological University.